OpenStudy (anonymous):

Simplify: -28^0 -1? or 1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[(-x)^0 = (x)^0 = 1\]

OpenStudy (anonymous):

@waterineyes It's not (-x)^0 though.. It's -x^0

OpenStudy (anonymous):

its -1

jimthompson5910 (jim_thompson5910):

-28^0 = -(28^0) = -(1) = -1 So -28^0 = -1

OpenStudy (anonymous):

If it is then -1 is the answer...

OpenStudy (anonymous):

Thanks for all the feedback everyone! :)